# SAFR: Neuron Redistribution for Interpretability

### v1

@@ Abstract
Superposition refers to encoding representations of multiple features within a single neuron, which is common in transformers. This property allows neurons to combine and represent multiple features, enabling the model to capture intricate information and handle complex tasks. Despite promising performance, the model’s interpretability has been diminished. This paper presents a novel approach to enhance transformer interpretability by regularizing feature superposition. We introduce SAFR  [^1] SAFR: Superposition-Aware Feature Regularization, which simply applies regularizations to the loss function to promote monosemantic representations for important tokens while encouraging polysemanticity for correlated token pairs, where important tokens and correlated token pairs are identified via VMASK Chen and Ji (2020) and attention weights. With a transformer model on two classification tasks, SAFR improves interpretability without compromising prediction performance. Given an input to the model, SAFR provides an explanation by visualizing the neuron allocation and interaction within the MLP layers.
@@ 1 Introduction
Individual neurons in neural networks can represent multiple features from the input. This phenomenon, known as superposition, improves the model’s ability to capture intricate relationships between features Olah et al. (2020) and complicates the understanding of the underlying processes behind the model’s decision-making Elhage et al. (2022). Facing these challenges, recent research like sparse autoencoders (SAEs) Huben et al. (2024) artificially decomposes the representation space into a sparse vector space through auxiliary networks. However, while SAEs provide a method to interpret features through combinations of sparse activations, there is still a lack of sufficient research on controlling neuron distribution.
In this paper, we ask the question: Can we enhance model interpretability by explicitly controlling the distribution of features across neurons? An intuitive approach is to encourage the development of monosemantic neurons Elhage et al. (2022); Bricken et al. (2023); Wang et al. (2024) by regulating activations. However, focusing solely on monosemanticity may limit the model’s ability to capture feature interactions, potentially hindering overall performance.
To address this challenge, we propose a novel method called Superposition-Aware Feature Regularization (SAFR) to enhance model interpretability by strategically reallocating neurons through a modified loss function, approached from two perspectives. As illustrated in Figure 1, our framework incorporates regularization techniques aimed at promoting monosemantic representations for important tokens, while simultaneously fostering polysemanticity among correlated token pairs. The identification of important tokens is achieved via a variational inference network, adapted from VMASK Chen and Ji (2020). Additionally, correlated token pairs are identified based on attention weights to foster polysemantic representation.
We evaluate SAFR on the SST-2 Socher et al. (2013) and IMDB Maas et al. (2011) datasets to assess its effectiveness. When the top 30% of words identified by SAFR are removed, we observe a significant drop in test accuracy—18.24% on SST-2 and 27.04% on IMDB. As we incrementally remove additional tokens, the accuracy continues to decline in a consistent manner. To further substantiate these findings, we visualize the neuron allocation within the FFN layers of a transformer block. Our experimental results confirm that SAFR can efficiently reallocate feature distributions across neurons while preserving a reasonable degree of polysemanticity, significantly improving interpretability in transformers while maintaining performance at comparable levels.
@@ 2 Preliminaries
Given an input sequence of $n$ tokens $S=\langle s_{1},\ldots,s_{n}\rangle$ and a transformer $f(\cdot)$ with $L$ layers, let $h^{\ell}_{i}$ denote the hidden representation obtained at layer $\ell\in[1,\ldots,L]$ and position $i\in[1,\ldots,n]$. Following previous work  Scherlis et al. (2022); Elhage et al. (2022), we define interference, polysemanticity, and capacity as follows.
@@
The interference ($I$) measures the overlap or similarity between the two representations at the same layer:
$I_{i,j}^{\ell}=\left(h^{\ell}_{i}\cdot h^{\ell}_{j}\right)^{2}$ (1)
Interference quantifies how much the hidden representations interfere with or influence each other. Higher interference means more overlap, suggesting the tokens share representational dimensions.
@@
Here, $\hat{h}^{\ell}_{i}$ denotes the normalized representation of $h^{\ell}_{i}$ (excluding its own magnitude). A high polysemanticity value indicates that a single dimension is “polysemantic”, meaning it represents information from multiple tokens simultaneously.
@@
The capacity ($C$) quantifies how much of a hidden representation is dedicated to representing token $i$:
@@
where $0\leq C_{i}\leq 1$ and $1\leq C\leq D$, $D$ denotes the hidden dimension size. A higher capacity value indicates that the representation at position $i$ is more focused on its own token.
@@ 3 Methodology
SAFR aims to enhance the interpretability of models by strategically reallocating neurons using superposition regularization strategy.
SST-2 IMDB
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15 80.88 74.31 76.92 3.96
$\lambda_{Imp}=0,\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26 79.07 73.30 64.70 14.37
SAFR 72.96 75.3 54.72 18.24 78.23 74.52 51.19 27.04
Table 1: Evaluation on SST-2 and IMDB datasets, k = 30. $\text{Acc}_{S}$ denotes prediction accuracy for original test dataset. $\text{Acc}_{\tilde{S}^{(r)}}$ denotes prediction accuracy for dataset randomly drop kwords. $\text{Acc}_{\tilde{S}^{(k)}}$ denotes prediction accuracy for deleting kwords based on capacity. SAFR achieves a SRS score of 18.24 for SST-2 and 27.04 for IMDB, outperforming the baseline model, indicating an improvement on model interpretability.
The baseline is defined via original model in cross-entropy loss $\mathcal{L}_{\text{CE}}=-\frac{1}{M}\sum_{m=1}^{M}\sum_{g=1}^{G}y_{m}^{g}\log(\hat{y}_{m}^{g})$, where $M$ is the number of samples in the dataset and $G$ is the number of classes in text classification. To improve model efficiency and interpretability in a constrained space, we propose a two-part regularization: one promotes monosemantic representations for important tokens, while the other encourages polysemanticity for correlated token pairs to reduce redundancy.
@@
We apply the VMASK Chen and Ji (2020) between the embedding layer and the positional encoder to identify important tokens. VMASK Chen and Ji (2020) is a variational word mask layer that is inserted into a neural text classifier and trained with the model. It learns to limit the flow of globally irrelevant or noisy word-level feature information to subsequent network layers, thus forcing the model to focus on the important features for prediction. To encourage monosemanticity for important tokens, we introduce a regularization term $\mathcal{L}_{\text{Importance}}=\sum_{i=1}^{T}\sqrt{P_{i}^{V}/{E}}$, where $T$ is the maximum token length, $E$ represents the embedding dimension and $P_{i}^{V}$ denotes the polysemanticity for the hidden representation of the $i$-th token in the VMASK layer. This regularization term penalize important tokens with high polysemanticity.
@@
We leverage the attention mechanism and employe the attention weights to identify correlated tokens. The $\alpha$-th self-attention head is described as follows:
@@
where $Q_{\alpha}=\mathbf{X}W^{Q}_{\alpha}$ and $\quad K_{\alpha}=\mathbf{X}W^{K}_{\alpha}$, with $Q_{\alpha},K_{\alpha}\in\mathbb{R}^{n\times d_{k}}$. $\mathbf{X}=E_{\text{pos}}(S^{\prime})$ denotes the input matrix to the attention layer, where $E_{\text{pos}}(S^{\prime})$ is the positional encoding applied to the output $S^{\prime}$ from the vmask layer. The score $A_{\alpha(i,j)}$ indicates how much attention token $i$ places on token $j$.
To encourage pairs of highly correlated features to exhibit high polysemanticity, we introduce a loss term $\mathcal{L}_{\text{Interaction}}=\sum_{\alpha}\sum_{i,j}\frac{1}{T^{2}}A_{\alpha(i,j)}\cdot(1-I_{i,j}^{A_{\alpha}})$, where $I_{i,j}^{A_{\alpha}}$ is the $Interference$ of the attention weights metric of the $\alpha$-th attention head. This loss term penalizes highly correlated features that exhibit low interference values.
@@
$\!=\!-\!\sum_{m=1}^{M}\!\sum_{g=1}^{G}y_{m}^{g}\!\log(\hat{y}_{m}^{g})\!+\!\lambda_{Imp}\!\cdot\!\sum_{m=1}^{M}\!\sum_{i=1}^{T}\!\sqrt{P_{i}^{V}\!/\!E}$
$\quad+\lambda_{Inter}\cdot\sum_{m=1}^{M}\sum_{\alpha}\sum_{i,j}\frac{1}{T^{2}}A_{\alpha(i,j)}\!(1-I_{i,j}^{A_{\alpha}})$
4 Experiment Setup
@@
We adopt 2 benchmark datasets: Stanford Sentiment Treebank binary version SST-2 Socher et al. (2013) and movie reviews IMDB Maas et al. (2011). Table 4 in Appendix A shows the statistics of the datasets.
@@
We use a typical transformer architecture with a single layer, following the standard setup Vaswani (2017). This includes the complete transformer framework with its attention mechanism and positional encodings. In the multi-layer perceptron (MLP) section, we use two feed forward sub-layers: we first expand the dimensionality by a factor of four, then reduce it back by the same factor, aligning with the commonly used configuration in transformer models. The model is using random embedding to avoid the influence of embedding information. Table 5 in Appendix A shows the statistics of the model.
@@
$SRS(k)\!=\!\frac{1}{M}\!(\sum_{S=1}^{M}\!1\!\cdot\!(\hat{y}_{S}=y)\!-\!\sum_{S=1}^{M}\!1\!\cdot\!(\hat{y}_{\tilde{S}^{(k)}}\!=\!y))$
where $\tilde{S}^{(k)}$ is constructed by dropping the $k\%$ top-scored words from $S$. The SRS metric serves to evaluate how effectively the model arranges neurons to represent features. The SRS metric quantifies the alignment between neuron allocations and the semantic significance of tokens. By assessing the structure of neuron allocation, SRS provides insights into the interpretability of the model’s internal encoding of information. Higher SRS values indicate that the deleted words are important, therefore indicates a stronger superposition regularization.
@@ 5 Results and Analysis
Results in Table 1 demonstrate that SAFR achieves superior SRS scores of 18.24 and 27.04 for SST-2 and IMDB datasets respectively, surpassing the baseline model. This provides evidence for SAFR’s efficacy in improving interpretability. Notably, even without regularization ($\lambda_{Imp}=0$ and $\lambda_{Inter}=0$), the interpretability layer before the transformer block yields improved SRS scores, suggesting its influence on neuronal allocation. An ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$ is detailed in Table 2. When calibrating these regularization terms, two critical trade-offs emerge: Excessive emphasis $\lambda_{Imp}$ may compromise the model’s ability to capture feature interactions. Conversely, focusing solely on $\lambda_{Inter}$ without balancing feature importance can reduce attention to critical words. Additional ablation studies can be found in Appendix D.
@@
Model $\text{Acc}_{S}$ $\text{Acc}_{\tilde{S}^{(r)}}$ $\text{Acc}_{\tilde{S}^{(k)}}$ SRS
Baseline 71.81 75.81 74.96 -3.15
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 71.24 71.98 61.98 9.26
$\lambda_{Imp}=0$, $\lambda_{Inter}=1$ 71.98 69.07 61.86 10.12
$\lambda_{Imp}=1$, $\lambda_{Inter}=0$ 71.47 72.16 55.35 16.12
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 72.96 75.30 54.72 18.24
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 61.46 62.04 51.06 10.40
Table 2: Ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$. $\lambda$ choosing influence prediction accuracy and model interpretability.
(a) Capacity Barchart of FFN1.
(b) Visualization of FFN1.
(c) Interference Heatmap of FFN1.
Figure 2: (a) Important words exhibit higher capacity. (b) Circle size represents feature capacity, with larger circles indicating greater capacity. Red lines denote positive correlations, blue lines indicate negative correlations, and shorter lines indicate stronger correlations. (c) Important words demonstrate lower polysemanticity, while correlated words pair exhibit relatively higher interference.
@@
Figure 3 illustrates the effect of token removal based on capacity. The ktop-scored words are directly removed from the original text to ensure that the evaluation reflects the model’s ability to perform with reduced input information. As tokens are incrementally removed, the accuracy declines consistently. The greater decline observed in our model compared to the baseline suggests that our model is more interpretable.
Figure 3: Sensitivity to $k$ Selection. As tokens are incrementally removed, the accuracy declines consistently.
@@
Table 3 presents an analysis of the average capacity per token status (important or not w.r.t VMASK) on the entire test dataset. Tokens are classified based on VMASK scores, with the top 30% identified as important and the rest 70% as less important. These results demonstrate the effectiveness of SAFR in prioritizing and allocating greater representational capacity to features that are important for tasks.
@@
All Tokens 0.1925 0.1372
Important Tokens 0.4141 0.2912
Less Important Tokens 0.0975 0.0712
@@
This strategic allocation enhances the model’s ability to provide clearer and more meaningful insights into its decision-making processes, reinforcing its value as a tool for model interpretability.
@@
We conducted an analysis of neuron allocation across layers, with the visualizations presented in Appendix C. Our observations reveal that neurons within the embedding layer exhibit a random distribution, failing to effectively capture their relative importance or inter-neuronal interactions. The VMASK layer begins to identify important tokens at a global level, yet it lacks the capability to analyze token-to-token interactions. While the attention layer demonstrates proficiency in capturing inter-token relationships, the interpretation of token importance remains a debatable topic. The second feedforward neural network layer compresses information into a lower-dimensional space and tend to allocate neuron capacity in a uniform manner. The expansion of the feedforward neural network, denotes FFN1 layer in our model, facilitates the network’s ability to learn more intricate patterns and relationships within the data Elhage et al. (2021). Our oberservation reveals that important tokens exhibit greater monosemanticity, and correlated token pairs demonstrate higher interference, illustrated in Figure 2(a) and 2(c).
Visualization of FFN1
Figure 2(b) illustrates the enhanced interpretability achieved after applying SAFR. In the visualization, the size of the circles represents the capacity of individual features, where larger circles indicate greater capacity. The red and blue lines depict interference between features, with red lines corresponding to positive correlations and blue lines indicating negative correlations. The length of these lines reflects the strength of the correlation. This representation provides a clear and intuitive understanding of both the distribution of neuron and the relationships among features, showcasing the effectiveness of SAFR in improving the interpretability of FFN1 layer. By visualizing these dynamics, the figure highlights key insights into feature importance and their interactions within the model.

### v2

@@ Abstract
Superposition refers to encoding representations of multiple features within a single neuron, which is common in deep neural networks. This property allows neurons to combine and represent multiple features, enabling the model to capture intricate information and handle complex tasks. Despite promising performance, the model’s interpretability has been diminished. This paper presents a novel approach to enhance model interpretability by regularizing feature superposition. We introduce SAFR, [^1] SAFR: Superposition-Aware Feature Regularization which simply applies regularizations to the loss function to promote monosemantic representations for important tokens while encouraging polysemanticity for correlated token pairs, where important tokens and correlated token pairs are identified via VMASK Chen and Ji (2020) and attention weights respectively. We evaluate SAFR with a transformer model on two classification tasks. Experiments demonstrate the effectiveness of SAFR in improving model interpretability without compromising prediction performance. Besides, SAFR provides explanations by visualizing the neuron allocation within the intermediate layers. [^2] The code can be found in https://github.com/chili-lab/SAFR.
@@ 1 Introduction
Individual neurons in neural networks can represent multiple features from the input. This phenomenon, known as superposition, improves the model’s ability to capture intricate relationships between features Olah et al. (2020), while also complicating the understanding of the underlying processes behind the model’s decision-making Elhage et al. (2022). Facing these challenges, recent research like sparse autoencoders (SAEs) Huben et al. (2024) artificially decomposes the activation space into a sparse vector space through auxiliary networks. While SAEs provide a method to interpret features through combinations of sparse activations, there is still a lack of sufficient research on controlling neuron distribution for interpretability.
In this paper, we ask the question: Can we enhance model interpretability by explicitly controlling the distribution of features across neurons? An intuitive approach is to encourage monosemantic neurons by regulating activations Elhage et al. (2022); Bricken et al. (2023); Wang et al. (2024). However, focusing solely on monosemanticity may limit the model’s ability to capture feature interactions, potentially hindering overall performance.
To address this challenge, we propose a novel method called Superposition-Aware Feature Regularization (SAFR) to enhance model interpretability by strategically redistributing neurons through a modified loss function, approached from two perspectives. As illustrated in Figure 1, our framework incorporates regularization techniques aimed at promoting monosemantic representations for important tokens, while simultaneously fostering polysemanticity among correlated token pairs. The identification of important tokens is achieved via a variational inference network, adapted from VMASK Chen and Ji (2020). Additionally, correlated token pairs are identified based on attention weights to foster polysemantic representation.
We evaluate the effectiveness of SAFR with a transformer model on the SST-2 Socher et al. (2013) and IMDB Maas et al. (2011) datasets. When the top 30% of words identified by SAFR are removed, we observe a significant drop in test accuracy—17.21% on SST-2 and 28.48% on IMDB. As we gradually remove additional words, accuracy continues to decline in a consistent manner. To further substantiate these findings, we visualize the neuron allocation within the FFN layers of a transformer block. Our experimental results demonstrate that SAFR can effectively redistribute features across neurons while preserving a reasonable degree of polysemanticity, significantly improving interpretability while maintaining prediction performance at comparable levels.
@@ 2 Preliminaries
Given an input sequence of $T$ tokens $S=\langle s_{1},\ldots,s_{T}\rangle$ and a neural network model $f(\cdot)$ with $L$ layers, let $h^{\ell}_{i}$ denote the hidden representation obtained at layer $\ell\in[1,\ldots,L]$ and token position $i\in[1,\ldots,T]$. Following previous work  Scherlis et al. (2022); Elhage et al. (2022), which defines interference, polysemanticity, and capacity based on features, we extend these definitions to hidden representations in our analysis.
@@
The interference ($I$) measures the overlap or similarity between two representations at the same layer:
$I_{i,j}^{\ell}=h^{\ell}_{i}\cdot h^{\ell}_{j}$ (1)
Interference quantifies how much the hidden representations interfere with each other. Higher interference indicates greater overlap, suggesting the tokens share representational dimensions.
@@
Here, $\hat{h}^{\ell}_{i}$ denotes the normalized representation of $h^{\ell}_{i}$ (divided by its own magnitude). A high polysemanticity value indicates that a single direction in the representation space is “polysemantic”, meaning it simultaneously represents information from multiple tokens.
@@
The capacity ($C$) quantifies how much of a hidden representation direction is dedicated to representing token $i$:
@@
where $0\leq C_{i}^{\ell}\leq 1$ and $1\leq C^{\ell}\leq T$ with $C^{\ell}=\sum_{i=1}^{T}C_{i}^{\ell}$ and $T$ denotes sequence length. A higher capacity value indicates that the representation at position $i$ is more focused on representing the $i$-th token in the input.
SST-2 IMDB
Model $\text{Acc}_{S}(\%)$ $\text{Acc}_{\tilde{S}^{(r)}}(\%)$ $\text{Acc}_{\tilde{S}^{(k)}}(\%)$ SRS $\text{Acc}_{S}(\%)$ $\text{Acc}_{\tilde{S}^{(r)}}(\%)$ $\text{Acc}_{\tilde{S}^{(k)}}(\%)$ SRS
Baseline 70.21 67.12 66.21 4.00 80.14 77.11 76.60 3.54
$\lambda_{Imp}=0,\lambda_{Inter}=0$ 72.56 69.47 64.44 8.12 78.43 76.10 73.56 4.87
SAFR 72.96 70.61 55.75 17.21 78.45 75.05 49.97 28.48
Table 1: Evaluation on SST-2 and IMDB datasets, k = 30. $\text{Acc}_{S}$ denotes prediction accuracy for original test dataset. $\text{Acc}_{\tilde{S}^{(r)}}$ and $\text{Acc}_{\tilde{S}^{(k)}}$ denote prediction accuracy after randomly deleting $k\%$ of tokens and deleting $k\%$ based on capacity, respectively. SAFR achieves a SRS score of 17.21 for SST-2 and 28.48 for IMDB, outperforming the baseline model, indicating an improvement on model interpretability. The optimal parameter settings for SAFR are $\lambda_{Imp}=0.1,\lambda_{Inter}=0.1$ for SST-2 and $\lambda_{Imp}=0.1,\lambda_{Inter}=1$ for IMDB.
@@ 3 Methodology
SAFR enhances model interpretability by strategically redistributing neurons through a superposition regularization strategy. By promoting a well-structured neuron distribution that balances importance and interaction, it makes token representations more meaningful.
The baseline is defined using the original model with cross-entropy loss $\mathcal{L}_{\mathrm{CE}}=-\frac{1}{N}\sum_{n=1}^{N}\sum_{g=1}^{G}y_{n}^{g}\log p_{n}^{g}$ , where $N$ is the number of samples in the dataset, and $G$ is the number of classes in text classification, $y_{n}^{g}$ is an indicator that equals 1 if sample $n$ belongs to class $g$ (and 0 otherwise), $p_{n}^{g}$ is the predicted probability for sample $n$ being of class $g$. To improve model interpretability in a constrained representational space, we propose a two-part regularization strategy: one that promotes monosemantic representations for important tokens, and the other encourages polysemanticity for correlated token pairs. This approach enables the model to effectively allocate its representational resources, as shown in Figure 1.
@@
We apply the VMASK Chen and Ji (2020) between the embedding layer and the positional encoder to select important tokens. A detailed introduction to VMASK is provided in Appendix A. To encourage monosemanticity for important tokens, we introduce a regularization term $\mathcal{L}_{\text{Importance}}=\frac{1}{T}\sum_{i=1}^{T}\sqrt{P_{i}^{V}/{E}}$, where $E$ represents the embedding dimension and $P_{i}^{V}$ denotes the polysemanticity for the hidden representation of the $i$-th token after the VMASK layer. This regularization penalizes important tokens with high polysemanticity.
@@
We leverage the attention mechanism and employ the attention weights to identify correlated tokens. The $\alpha$-th self-attention head is described as follows:
@@
where $\alpha\in[1,\ldots,M]$, $Q_{\alpha}=\mathbf{X}W^{Q}_{\alpha}$ and $\quad K_{\alpha}=\mathbf{X}W^{K}_{\alpha}$, with $Q_{\alpha},K_{\alpha}\in\mathbb{R}^{T\times d_{k}}$, $d_{k}$ denotes the dimension of the key and query vectors. $\mathbf{X}=E_{\text{pos}}(S^{\prime})\in\mathbb{R}^{T\times E}$ denotes the input matrix to the attention layer, where $E_{\text{pos}}(S^{\prime})$ is the positional encoding applied to the output $S^{\prime}$ from the VMASK layer. The score $A_{\alpha(i,j)}$ indicates how much attention token $i$ places on token $j$.
To encourage highly correlated token pairs to exhibit high polysemanticity, we introduce a loss term $\mathcal{L}_{\text{Interaction}}=\sum_{\alpha}\sum_{i,j}\frac{1}{T^{2}}A_{\alpha(i,j)}\cdot(1-I_{i,j}^{A_{\alpha}})$, where $I_{i,j}^{A_{\alpha}}$ is the $Interference$ of the attention weights matrix for the $\alpha$-th attention head. This loss term penalizes highly correlated tokens that exhibit low interference values.
@@
$\mathcal{L}$ $=\mathcal{L}_{\text{CE}}+\lambda_{Imp}\cdot\mathcal{L}_{\text{Imp}}+\lambda_{Inter}\cdot\mathcal{L}_{\text{Inter}}$
$=\mathcal{L}_{\text{CE}}+\lambda_{Imp}\!\cdot\!\frac{1}{NT}\sum_{n=1}^{N}\sum_{i=1}^{T}\sqrt{\frac{P_{i}^{V}}{E}}$
$\quad+\lambda_{Inter}\!\cdot\!\frac{1}{NMT^{2}}\!\sum_{n=1}^{N}\!\sum_{\alpha=1}^{M}\!\sum_{i,j}\!A_{\alpha(i,j)}\!(1\!-\!I_{i,j}^{A_{\alpha}})$
4 Experimental Setup
@@
We adopt two benchmark datasets: Stanford Sentiment Treebank binary version SST-2 Socher et al. (2013) and movie reviews IMDB Maas et al. (2011). Table 4 in Appendix B presents the dataset statistics.
@@
We use a typical transformer architecture with a single layer, following the standard setup Vaswani (2017). This includes the complete transformer framework with its attention mechanism and positional encoding. In the multi-layer perceptron (MLP) section, we use two fully connected layers: we first expand the dimensionality by a factor of four, apply a ReLU activation, then reduce it back by the same factor, aligning with the commonly used configuration in transformer models. The model uses random embeddings to avoid the influence of pre-trained embedding information. Table 5 in Appendix B presents the model statistics.
@@
$SRS(k)\!=\!\frac{1}{N}\!(\sum_{S=1}^{N}\!1\!\cdot\!(\hat{y}_{S}=y)\!-\!\sum_{S=1}^{N}\!1\!\cdot\!(\hat{y}_{\tilde{S}^{(k)}}\!=\!y))$
where $\tilde{S}^{(k)}$ is constructed by dropping the $k\%$ top-scored tokens identified by SAFR from $S$. The SRS measures how effectively the model arranges neurons. The SRS metric quantifies the alignment between neuron allocations and the semantic significance of tokens. By assessing the structure of neuron allocation, SRS provides insights into the interpretability of the model’s internal encoding of information. Higher SRS values indicate that the removed words were highly important, signifying stronger superposition regularization.
(a) Capacity Barchart of FC1.
(b) Visualization of FC1.
(c) Interference Heatmap of FC1.
Figure 2: (a) Important tokens exhibit higher capacity. (b) Circle size represents capacity, with larger circles indicating greater capacity. Red lines denote positive correlations, blue lines indicate negative correlations, and shorter lines indicate stronger correlations. (c) Important tokens demonstrate lower polysemanticity, while correlated token pair exhibit relatively higher interference.
@@ 5 Results and Analysis
After applying SAFR, the first fully connected layer redistributes neurons, making Capacity interpretable by jointly capturing interference and polysemanticity, thereby reflecting both interaction and importance, as validated by SRS evaluation. Table 1 shows that SAFR achieves SRS scores of 17.21 and 28.48 on the SST-2 and IMDB datasets, outperforming the baseline and demonstrating improved interpretability. Notably, even without regularization ($\lambda_{Imp}=0$, $\lambda_{Inter}=0$), the VMASK layer before the Transformer block
@@
Model $\text{Acc}_{S}(\%)$ $\text{Acc}_{\tilde{S}^{(r)}}(\%)$ $\text{Acc}_{\tilde{S}^{(k)}}(\%)$ SRS
Baseline 70.21 67.12 66.21 4.00
$\lambda_{Imp}=0$, $\lambda_{Inter}=0$ 72.56 69.47 64.44 8.12
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=0.1$ 72.96 70.61 55.75 17.21
$\lambda_{Imp}=0.1$, $\lambda_{Inter}=1$ 72.67 69.18 59.58 13.09
$\lambda_{Imp}=1$, $\lambda_{Inter}=0.1$ 71.81 69.13 56.38 15.43
$\lambda_{Imp}=100$, $\lambda_{Inter}=100$ 63.01 61.64 54.20 8.81
Table 2: Ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$. The choice of $\lambda$ affects both prediction accuracy and model interpretability.
enhances SRS scores, indicating its positive effect on neuronal allocation. Table 2 presents an ablation study of $\lambda_{Imp}$ and $\lambda_{Inter}$, highlighting two trade-offs: excessive $\lambda_{Imp}$ may compromise the model’s ability to capture feature interactions, while high $\lambda_{Inter}$ without balancing feature importance can reduce attention to critical tokens. Additional ablation studies are in Appendix D.
@@
Figure 3 illustrates the effect of token removal based on capacity. The ktop-scored tokens are directly removed from the original text to ensure that the evaluation reflects the model’s ability to perform with reduced input information. As tokens are gradually removed, accuracy declines consistently. The greater decline in our model compared to the baselines suggests better interpretability.
Figure 3: Sensitivity to $k$ Selection. As tokens are gradually removed, accuracy declines consistently.
@@
Table 3 presents an analysis of the average capacity per token status (important or not w.r.t VMASK) on the entire test set. Tokens are categorized based on VMASK scores, with the top 30% identified as important and the rest 70% as less important. These results demonstrate SAFR’s effectiveness in prioritizing and allocating greater representational capacity to task-relevant tokens. This strategic allocation improves the model’s clarity in decision-making.
@@
All Tokens 0.2981 0.1403
Important Tokens (top 30%) 0.5745 0.2035
Less Important Tokens (the rest 70%) 0.1794 0.1132
@@
We analyzed neuron allocation across layers, with visualizations in Appendix C. Neurons in the embedding layer are randomly distributed, failing to effectively capture their relative importance or inter-neuronal interactions. The VMASK layer begins to identify important tokens at a global level, yet it lacks the capability to analyze token-to-token interactions. While the attention layer demonstrates proficiency in capturing inter-token relationships, the interpretation of token importance remains debated Jain and Wallace (2019). The second fully connected layer compresses information into a lower-dimensional space and tend to allocate neuron capacity in a uniform manner. The expansion step in the feedforward network (FC1 layer in our model) allows the network to capture high-dimensional representations of the data. Our observation of this layer reveals that important tokens exhibit greater monosemanticity, and correlated token pairs demonstrate higher interference, illustrated in Figure 2(a) and 2(c).
Visualization of FC1
Figure 2(b) illustrates the enhanced interpretability achieved after applying SAFR. In the visualization, the size of the circles represents the capacity of individual tokens, where larger circles indicate greater capacity. The red and blue lines depict interference between tokens, with red lines corresponding to positive correlations and blue lines indicating negative correlations. The length of these lines reflects the strength of the correlation. This visualization illustrates neuron distribution and token relationships, demonstrating SAFR’s effectiveness in enhancing feedforward layer interpretability. By visualizing these dynamics, the figure highlights key insights into token importance and their interactions within the model.